OpenStudy (anonymous):

Question: \[\mathcal{L}\left \{ (3t+1)\mathcal{U}(t-1) \right \}\]Properties: \[\mathcal{L}\left \{ f(t-a)\mathcal{U}(t-a) \right \}=e^{-as}F(s)\]Procedure: We see that\[f(t)=3t+1\]\[f(t-1)=3(t-1)+1=3t-3+1=3t-2.\]Therefore,\[\mathcal{L}\left \{ (3t-2+3)\mathcal{U}(t-1) \right \}=\mathcal{L}\left \{ (3t-2)\mathcal{U}(t-1) \right \}+\mathcal{L}\left \{ 3\mathcal{U}(t-1) \right \}\implies\]\[e^{-as}\left (\frac{3}{s^2}-\frac{2}{s}\right )+3\frac{e^{-as}}{s}\]Does this seem plausible?

OpenStudy (unklerhaukus):

yerp

OpenStudy (anonymous):

I guess this could be further simplified to:\[e^{-s}\left ( \frac{3}{s^2}+\frac{1}{s} \right ).\]

OpenStudy (unklerhaukus):

dont forget the "a" in the exponent

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/Classes/DE/Laplace_Table.aspx Look at number 28 on the table.

OpenStudy (anonymous):

Thank you so much, Joe.